OpenStudy (anonymous):

one company calculates the yearly cost of insuring a new car with the formula show below. [ C= cost, a= age of owner, p= price of the car ] C = 0.0006p[140 - 1.5a] How much will it cost a 35 year old driver to insure a new $24,000 car?

OpenStudy (jim766):

C = 0.0006p(140 - 1.5a) C = 0.0006(24,000)(140 - 1.5(35)) C = 14.4(140-52.5) C = 14.4(87.5) C = $1,260